OpenStudy (anonymous):

If f(x) = x + 2, find f(3). A.) 1 B.) 6 C.) 5 D.) -1

OpenStudy (shamil98):

substitute 3 for x

OpenStudy (wolfe8):

Turtle, do you want an explanation?

OpenStudy (anonymous):

so, the next step is f(2) x 4...right?

OpenStudy (shamil98):

f(x) = x + 2 if f (3) you put 3 into x..

OpenStudy (anonymous):

srry...f(2)=4

OpenStudy (shamil98):

yes. 2+2 = 4

OpenStudy (anonymous):

wow...im not THAT stupid

OpenStudy (shamil98):

I never said you were ._. im saying if you plug 2 into f(x) = x + 2 that would be your answer, confirming it ..

OpenStudy (anonymous):

so would it be B?

OpenStudy (shamil98):

f(x) = x + 2 f(3) = 3 + 2

OpenStudy (anonymous):

srry...i ment to write C...is it C?

OpenStudy (shamil98):

Yes.

OpenStudy (anonymous):

thx